Shift report 20020121 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 21 2002 - 16:21:48 EST

  • Next message: Flemming Videbaek: "Fw: Workshop on Glueball and Central Production at RHIC"

    ------------------------------------------------------------
    Date: Mon Jan 21 16:21:47 2002, Shift: 08:00-16:00
    Supervisor: d Beavis, with: 
    Log book#12 pages 99 - 100
    
    Shift Summary:
    ------------------------------------------------------------
    Morning was with a poor store.
    Took one data run.  singles to coin. on inelastic counters was between 0.5 and 0.3 so probably poor eff.
    
    Run 6723 had 18k trigger 3 and 16 k of trigger 6 with BFS
    
    Changed D5 to B polarity
    
    Had access to IR with C. Pearson and was able to get BFS to move after removing motor and then placing it back on.
    
    2 Hours of beam this shift.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 21 2002 - 16:22:42 EST